CAN YOU PLEASE answer my question correctly, i posted this question and got a wrong answer. also show how you got the answer, THANK YOU. Given the following characteristics of magnetic tape, and using linear recording: - Density=2000 bpi - Speed= 1500 inches/second - Size=3000 feet - Number of records to be stored: 200,000 records - Size of each records=100 bytes - Block size=10 records - IBG=.5 inches Calculate the following: 1- Number of blocks needed 2- Size of the block in bytes 3- Time required to read one block 4- Time required to read all blocks 5- Amount of tape used for data only in inches 6- Total amount of tape used (data + IBG) in inches
Answer #1

1. Blocks needed = Total number of records / number of record per block = 200,000 / 10 = 20,000 blocks

2. Size of block in bytes = Size of record * number of records per block =100 bytes/rec * 10 rec/block = 1000 bytes/block

3. Time to read one block:

  • data time = 1000 bytes/block / 2000 bytes/in / 1500 in/sec = 0.00033 sec/block
  • ibg time = 0.5 in/block / 1500 in/sec = 0.00033 sec/block
  • total = 0.00033 + 0.00033 = 0.00066 sec/block


4. Time to read all blocks = time to read 1 block * total number of blocks = 200,000 blocks * 0.00066 sec/block = 133.33 sec

5. Data length = 200,000 blocks * 0.00033 sec/block * 1500 in/sec = 99,000 in

6. Total amount of tape used = 133.33 sec * 1500 in/sec = 199995 in